A 9‐year 6.5‐percent coupon bond is selling for 104.7 percent of par. What is the bond’s market yield if it makes semi‐annual coupon payments?
Answer #1

Suppose the face value of the bond be $1000
Given that the bond is selling for 104.7 percent of par. It means the current or present value is 1.047 times the face value=1000*1.04=1047

Coupon rate=6.5%
Semiannual coupon rate=6.5%/2
Semiannual coupon payment=(Semiannual coupon rate)*(Face value)=(0.0325)*(1000)=32.5

Time period is 9 years
As the coupon payments are made semiannually, the number of periods=9*2=18

We can calculate the yield using excel.

U V 1 Face value 1000 2 Present value -1047 Number of period 3 18 4 Payment 32.5 5 Yield 2.91% 6 Formula used: RATE(V3,V4,V2,

As the present value is a cash outflow, we have taken it as negative in excel.
The semiannual yield is 2.91%
So, the annual yield is 2.91%*2=5.82%